# Learning to Partially Defer for Sequences

### v1

@@ Appendix D Training Details and Hyperparameters
Task/Hyperparameters TSP-OneTime XSUM-Tokenlevel XSUM-OneTime
Learning rate 5e-4 1e-3 5e-4
Early Stopping (patience, delta) 20, 1e-4 7, 1e-4 20, 1e-4
Architecture MLP LSTM MLP
Hidden layers 1 3 1
Hidden units 8 128 8
Dropout 0.2 0.4 0.2
Epochs 200 100 200
Training Samples 2400 800 2400
Gradient Clipping Norm 1.0 1.0 1.0
Weight Decay 0.005 0.001 0.005
@@
TSP: Using the RL4CO library [56], TSP graphs with $50$ nodes were generated by sampling the node coordinates from a standard normal distribution. These samples were fed into a PointerNet (predictor) trained with $10000$ samples. The node embeddings, graph context, and log scores from the PointerNet served as inputs for the One-Time rejector. This rejector only took a maximum of $10$ minutes to train.
XSUM: During training and inference, Tokenlevel rejector was only exposed to samples with summaries of less than $15$ words. The model was trained without teacher forcing. It took about $5$ hours to train. The token level rejector $r_{1}$ recieves the encoder hidden states from the predictor as hidden states and the decoder hidden states as input. This continues in an autoregressive fashion. We defer the reader to Appendix G for ablation studies on various training strategies for the tokenlevel rejector.
@@
Appendix E Effects of $\alpha_{j}$
@@
Appendix F AUDC and Percent Improvement for Size of ${\cal J}$ experiment
@@
Appendix G Token-level Rejector Ablation
${\cal L}^{\text{Token}}$ bears resemblance to a weighted multi-label classification loss function. Specifically, the token-level rejector follows a classifier chain approach [57]. Classifier chains is a multi-label classification (MLC) approach that involves training $L$ binary classifiers in a pre-specified order with the $j^{\text{th}}$ binary classifier taking the feature vector and the previous $j-1$ labels as input, forming a chain of predictors. At inference time, the test features are fed into the chain such that the previous label predictions are appended to the features for the next label classifier. Our token-level rejector $r_{j}$ implicitly makes these chained predictions by taking in $\widehat{y}_{<j}$ which is informed by the previous $(j-1)$ rejectors’ rejection decision.
With this comparison established, model architectures and training techniques from classifier chain literature can be utilized. Recurrent classifier chains [58] have emerged as a popular choice for MLC as they efficiently maximize accuracy without requiring the large number of trainable parameters typically needed for $K$ separate binary classifiers.
Since the inputs of recurrent classifier chains at training time are typically ground truth labels, training token-level rejectors would similarly require using the correct rejection decision, $\mathbbm{1}_{l\left(y,\widehat{y}_{j}^{h}\right)>c_{j}(x,\widehat{y}_{<j},y)}$, to determine the $j^{\text{th}}$ token prediction for the next label rejector. Following this strategy, also called “teacher forcing” [59], can often cause a distribution shift, as label predictions will be used at test time instead of the correct label, resulting in errors propagating through the chain. Scheduled sampling [27] offers a middle ground by labeling the training data with the correct rejection decision with probability $p$, and otherwise using the predicted one. This probability $p$ often decays at a linear or exponential rate over epochs, gradually reducing the model’s dependence on ground-truth rejection rules.

### v2

@@ Appendix D Training Details and Hyperparameters
Task/Hyperparameters TSP-OneTime XSUM-Tokenlevel XSUM-OneTime MWP-Tokenlevel MWP-OneTime
Learning rate 5e-4 1e-3 5e-4 5e-4 5e-4
Early Stopping (patience, delta) 20, 1e-4 7, 1e-4 20, 1e-4 7, 1e-4 20, 1e-4
Architecture MLP LSTM MLP LSTM MLP
Hidden layers 1 3 1 2 1
Hidden units 8 128 8 64 8
Dropout 0.2 0.4 0.2 0.4 0.2
Epochs 200 100 200 100 200
Training Samples 2400 800 2400 70080 70080
Gradient Clipping Norm 1.0 1.0 1.0 1.0 1.0
Weight Decay 0.005 0.001 0.005 0.001 0.005
@@
TSP: Using the RL4CO library Berto et al. [2023], TSP graphs with $50$ nodes were generated by sampling the node coordinates from a standard normal distribution. These samples were fed into a PointerNet (predictor) trained with $10000$ samples. The node embeddings, graph context, and log scores from the PointerNet served as inputs for the One-Time rejector. This rejector only took a maximum of $10$ minutes to train.
XSUM: During training and inference, Tokenlevel rejector was only exposed to samples with summaries of less than $15$ words. The model was trained without teacher forcing. It took about $5$ hours to train. The token level rejector $r$ recieves the encoder hidden states from the predictor as hidden states and the decoder hidden states as input. This continues in an autoregressive fashion. We defer the reader to Appendix H for ablation studies on various training strategies for the tokenlevel rejector.
@@
MWP: The token-level rejector took about $5$ hour to train. e token level rejector $r$ recieves the encoder hidden states from the predictor as hidden states and the decoder hidden states as input. This continues in an autoregressive fashion. The OneTime rejector was trained on a dataset containing all the varying mixtures of predictions which took a couple of minutes to generate. This rejector takes the encoder hidden states, decoder hidden states, and monte carlo variance estimates from the predictor as input.
Appendix E More Cost-Loss Plots
See Figure 3
Figure 3: Cost-Loss plots for TSP (on the left) and MWP (on the right).
Appendix F Effects of $\alpha_{j}$
@@
Appendix G AUDC and Percent Improvement for Size of $\cal J$} experiment
@@
Appendix H Token-level Rejector Ablation
${\cal L}^{\text{Token}}$ bears resemblance to a weighted multi-label classification loss function. Specifically, the token-level rejector follows a classifier chain approach Read et al. [2021]. Classifier chains is a multi-label classification (MLC) approach that involves training $L$ binary classifiers in a pre-specified order with the $j^{\text{th}}$ binary classifier taking the feature vector and the previous j1j-1 labels as input, forming a chain of predictors. At inference time, the test features are fed into the chain such that the previous label predictions are appended to the features for the next label classifier. Our token-level rejector $r_{j}$ implicitly makes these chained predictions by taking in $\widehat{y}_{<j}$ which is informed by the previous (j1)(j-1) rejectors’ rejection decision.
With this comparison established, model architectures and training techniques from classifier chain literature can be utilized. Recurrent classifier chains Nam et al. [2017] have emerged as a popular choice for MLC as they efficiently maximize accuracy without requiring the large number of trainable parameters typically needed for $K$ separate binary classifiers.
Since the inputs of recurrent classifier chains at training time are typically ground truth labels, training token-level rejectors would similarly require using the correct rejection decision, $\mathbbm{1}_{l\left(y,\widehat{y}_{j}^{h}\right)>c_{j}(x,\widehat{y}_{<j},y)}$, to determine the $j^{\text{th}}$ token prediction for the next label rejector. Following this strategy, also called “teacher forcing” Williams and Zipser [1989], can often cause a distribution shift, as label predictions will be used at test time instead of the correct label, resulting in errors propagating through the chain. Scheduled sampling Bengio et al. [2015] offers a middle ground by labeling the training data with the correct rejection decision with probability $p$, and otherwise using the predicted one. This probability $p$ often decays at a linear or exponential rate over epochs, gradually reducing the model’s dependence on ground-truth rejection rules.